# PoLi-RL: A Point-to-List Reinforcement Learning Framework for Conditional Semantic Textual Similarity

### v1

@@ 4.2 Qualitative Analysis: A Case Study on Nuanced Reasoning
To qualitatively analyze our framework’s nuanced reasoning ability, we present a case study in Figure 4 on the challenging condition “The person’s connection with the ground”. The baseline models show clear deficiencies: the Few-shot model exhibits a brittle, literal interpretation, fixating on the preposition ‘on’ while ignoring the real-world context, leading to an incorrect prediction. The SFT Baseline, while achieving a closer score, shows uncertainty and fails to make the necessary fine-grained distinction. In contrast, PoLi-RL applies world knowledge to infer that both “jumping” and “doing a flip” result in the same “not connected” state, arriving at the accurate score of 4 by identifying the differing actions leading to the state as unimportant details. This case study demonstrates that PoLi-RL cultivates the contextual inference ability required for complex conditional judgments.
@@ 5 Conclusion
We introduce PoLi-RL, a two-stage reinforcement learning framework that resolves the complex credit assignment problem associated with listwise objectives in C-STS through a progressive curriculum and an innovative Parallel-Slice Ranking Reward (PSRR) mechanism. Our method establishes a new state-of-the-art for the cross-encoder architecture, significantly outperforming SFT baselines and even large models like GPT-4. As the first successful application of reinforcement learning to this task, our study validates a powerful paradigm for aligning LLMs with complex ranking objectives, paving the way for future applications in retrieval and other ranking-based tasks.
6 Contact Information
Emails:    Zixin Song: songzx24@mails.tsinghua.edu.cn    Bowen Zhang: zbw23@mails.tsinghua.edu.cn
Qian-Wen Zhang: cowenzhang@tencent.com
@@ A.1 Related Work
Conditional Semantic Textual Similarity
@@
C-STS is a recent advancement over traditional STS that introduces a natural language condition to disambiguate the measurement of similarity between two texts. Research in this area has primarily established three mainstream paradigms: Bi-encoder, Tri-encoder, and Cross-encoder. Given two texts and a condition, the bi-encoder architecture typically uses a Siamese network to encode two text-condition pairs, while the tri-encoder architecture encodes each text and the condition separately before an aggregation step. A prevalent optimization strategy for these paradigms is contrastive learning. For instance, [4] propose a conditional contrastive learning framework that pulls representations of the same text pair under a high-similarity condition closer, while pushing them apart under a low-similarity one. Extending this, Hyper-CL [20] utilizes a tri-encoder setup where a hypernetwork generates condition-specific projectors to dynamically adapt sentence representations within a contrastive objective. More recently, CSR introduced a parameter-free router for the tri-encoder, using the condition to re-weight sentence tokens to amplify relevant information without increasing model size.
@@
In contrast, the cross-encoder architecture processes the concatenated texts and condition as a single input, enabling deep, token-level interaction. However, this theoretical advantage did not consistently translate to superior performance in earlier discriminative models. The state-of-the-art method in this setting, SEAVER, addressed this discrepancy by identifying that such models can be distracted by condition-irrelevant tokens. To resolve this, SEAVER [6] introduces an attention reallocation mechanism that optimizes the model by re-weighting internal attention scores during fine-tuning, forcing a focus on the most salient information.
@@
The advent of LLMs has introduced new approaches, mainly few-shot inference and using LLMs as feature extractors [7] [21]. However, all these prior works, regardless of architecture, are confined to supervised paradigms like SFT or contrastive learning. Our work is the first to apply reinforcement learning to this task.
Reinforcement Learning for Large Language Models
@@
Reinforcement Learning (RL) is pivotal for aligning and enhancing Large Language Models (LLMs), with algorithms like Proximal Policy Optimization (PPO) [22] widely used to optimize non-differentiable objectives in tasks such as reasoning and code generation. To address PPO’s limitations, such as high variance in long-sequence tasks, advanced variants have emerged.
@@
Group Relative Policy Optimization (GRPO) [8] addresses these limitations by introducing a group-relative advantage estimation, which eliminates the need for a separate value function through Z-score normalization of rewards within sample groups. This approach simplifies training and enhances sample efficiency, as demonstrated in models like DeepSeek-R1. Building on this, Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) [13] provides an open-source, scalable RL system tailored for LLMs. DAPO incorporates key improvements, including dynamic sampling to adaptively adjust the number of generated completions based on reward variance, making it particularly effective for long-horizon reasoning tasks.
@@ A.2 Prompt Template for PoLi-RL
Prompt for C-STS Task Judge the semantic similarity between Sentence 1 and Sentence 2 based completely on the given Condition. The final output must be exactly in this format: the similarity judgment (‘yes’ or ‘no’) followed by the score in parentheses, wrapped in <answer></answer> tags. Examples: <answer>yes(4)</answer>, <answer>no(1)</answer>. Include no other text, tags, or explanations. To arrive at this output, follow these two steps: Step 1: Binary Judgment. Determine if the sentences are ‘similar’ (‘yes’) or ‘not similar’ (‘no’). ‘similar’: The sentences are roughly, mostly, or completely equivalent under the condition. ‘not similar’: The sentences are dissimilar under the condition. Step 2: Fine-grained Score. Assign an integer score based on Step 1: For a ‘yes’ judgment: * 5: The two sentences are completely equivalent as they mean the same thing with respect to the condition. * 4: The two sentences are mostly equivalent, but some unimportant details differ with respect to the condition. * 3: The two sentences are roughly equivalent, but some important information differs or is missing with respect to the condition. For a ‘no’ judgment: * 2: The two sentences are dissimilar, but are on a similar topic with respect to the condition or shares a close semantic relationship. This applies when items are clearly different, but not direct opposites. * 1: The two sentences are dissimilar with respect to the condition, representing a direct opposition or a clear, unrelated difference. (e.g., ‘man’ vs. ‘woman’). Here are some examples: Example 1:
<Sentence1>: A girl is cooking in a kitchen and a man is standing next to her.
<Sentence2>: A man sitting with a pizza in his hand in front of pizza on the table.
<Condition>: The number of people.
<answer>no(1)</answer>
Explanation: The first sentence mentions two people, while the second sentence mentions only one person.
Example 2:
<Sentence1>: A wood table sitting by a wood framed bed with a lamp on it.
<Sentence2>: A microwave, refrigerator, television, and wooden drawers sit in the corner of a bedroom.
<Condition>: The room type.
<answer>yes(5)</answer>
Explanation: We can infer from the two sentences that the room type are both bedroom.
Example 3:
<Sentence1>: A small crowd gathered around the injured person.
<Sentence2>: A crowd jumps up and down to the tunes played by an artist.
<Condition>: The number of people
<answer>yes(3)</answer>
Explanation: While both sentences mention crowds, it is important and unclear how many people there are.
Now, apply these steps to the following sentences: <Sentence1>: {sentence1}
<Sentence2>: {sentence2}
<Condition>: {condition}

### v2

@@ 4.2 Qualitative Analysis: A Case Study on Nuanced Reasoning
To qualitatively analyze our framework’s nuanced reasoning ability, we present a case study in Figure 4 on the challenging condition “The person’s connection with the ground”. The baseline models show clear deficiencies: the few-shot model exhibits a brittle, literal interpretation, fixating on the preposition ‘on’ while ignoring the physical context, leading to an incorrect prediction. The SFT Baseline, while closer to the target, lacks the reasoning depth to make the necessary fine-grained distinction. In contrast, PoLi-RL applies world knowledge to infer that both “jumping” and “doing a flip” result in the same “not connected” state, arriving at the accurate score of 4 by identifying the differing actions leading to the state as unimportant details. This case study demonstrates that PoLi-RL cultivates the contextual inference ability required for complex conditional judgments.
@@ 5 Conclusion
We introduce PoLi-RL, a two-stage reinforcement learning framework that resolves the coarse-grained credit assignment problem associated with listwise objectives in C-STS through a progressive curriculum and an innovative Parallel-Slice Ranking Reward (PSRR) mechanism. Our method establishes a new state-of-the-art for the cross-encoder architecture, significantly outperforming SFT baselines and even proprietary reasoning models like GPT-4o. As the first successful application of reinforcement learning to this task, our study validates a powerful paradigm for aligning LLMs with complex ranking objectives, demonstrating the framework’s potential to other ranking-based tasks.
@@ A.1 Related Work
Conditional Semantic Textual Similarity.
@@
C-STS is a recent advancement over traditional STS (Nakov et al., 2013; Zhang et al., 2024a; 2025b) that introduces a natural language condition to disambiguate the measurement of similarity between two texts. Research in this area has primarily established three mainstream paradigms: Bi-encoder, Tri-encoder, and Cross-encoder. Given two texts and a condition, the bi-encoder architecture typically uses a Siamese network to encode two text-condition pairs, while the tri-encoder architecture encodes each text and the condition separately before an aggregation step. A prevalent optimization strategy for these paradigms is contrastive learning. For instance, (Liu et al., 2025) propose a conditional contrastive learning framework that pulls representations of the same text pair under a high-similarity condition closer, while pushing them apart under a low-similarity one. Extending this, Hyper-CL (Yoo et al., 2024) utilizes a tri-encoder setup where a hypernetwork generates condition-specific projectors to dynamically adapt sentence representations within a contrastive objective. More recently, CSR introduced a parameter-free router for the tri-encoder, using the condition to re-weight sentence tokens to amplify relevant information without increasing model size.
@@
In contrast, the cross-encoder architecture processes the concatenated texts and condition as a single input, enabling deep, token-level interaction. However, this theoretical advantage did not consistently translate to superior performance in earlier discriminative models. The state-of-the-art method in this setting, SEAVER, addressed this discrepancy by identifying that such models can be distracted by condition-irrelevant tokens. To resolve this, SEAVER (Li et al., 2024) introduces an attention reallocation mechanism that optimizes the model by re-weighting internal attention scores during fine-tuning, forcing a focus on the most salient information.
@@
The advent of LLMs has introduced new approaches, mainly few-shot inference and using LLMs as feature extractors (Yamada & Zhang, 2025) (Li & Li, 2023). However, all these prior works, regardless of architecture, are confined to supervised paradigms like SFT or contrastive learning. Our work is the first to apply reinforcement learning to this task.
Reinforcement Learning for Large Language Models.
@@
Reinforcement Learning (RL) is pivotal for aligning and enhancing Large Language Models (LLMs), with algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017) widely used to optimize non-differentiable objectives in tasks such as reasoning and code generation. To address PPO’s limitations, such as high variance in long-sequence tasks, advanced variants have emerged.
@@
Group Relative Policy Optimization (GRPO) (Guo et al., 2025) addresses these limitations by introducing a group-relative advantage estimation, which eliminates the need for a separate value function through Z-score normalization of rewards within sample groups. This approach simplifies training and enhances sample efficiency, as demonstrated in models like DeepSeek-R1. Building on this, Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025) provides an open-source, scalable RL system tailored for LLMs. DAPO incorporates key improvements, including dynamic sampling to adaptively adjust the number of generated completions based on reward variance, making it particularly effective for long-horizon reasoning tasks.
@@ A.2 Prompt Template for PoLi-RL
Prompt for C-STS Task Judge the semantic similarity between Sentence 1 and Sentence 2 based completely on the given Condition. The final output must be exactly in this format: the similarity judgment (‘yes’ or ‘no’) followed by the score in parentheses, wrapped in <answer></answer> tags. Examples: <answer>yes(4)</answer>, <answer>no(1)</answer>. Include no other text, tags, or explanations. To arrive at this output, follow these two steps: Step 1: Binary Judgment. Determine if the sentences are ‘similar’ (‘yes’) or ‘not similar’ (‘no’). ‘similar’: The sentences are roughly, mostly, or completely equivalent under the condition. ‘not similar’: The sentences are dissimilar under the condition. Step 2: Fine-grained Score. Assign an integer score based on Step 1: For a ‘yes’ judgment: * 5: The two sentences are completely equivalent as they mean the same thing with respect to the condition. * 4: The two sentences are mostly equivalent, but some unimportant details differ with respect to the condition. * 3: The two sentences are roughly equivalent, but some important information differs or is missing with respect to the condition. For a ‘no’ judgment: * 2: The two sentences are dissimilar, but are on a similar topic with respect to the condition or shares a close semantic relationship. This applies when items are clearly different, but not direct opposites. * 1: The two sentences are dissimilar with respect to the condition, representing a direct opposition or a clear, unrelated difference. (e.g., ‘man’ vs. ‘woman’). Here are some examples: Example 1: <Sentence1>: A girl is cooking in a kitchen and a man is standing next to her. <Sentence2>: A man sitting with a pizza in his hand in front of pizza on the table. <Condition>: The number of people. <answer>no(1)</answer> Explanation: The first sentence mentions two people, while the second sentence mentions only one person. Example 2: <Sentence1>: A wood table sitting by a wood framed bed with a lamp on it. <Sentence2>: A microwave, refrigerator, television, and wooden drawers sit in the corner of a bedroom. <Condition>: The room type. <answer>yes(5)</answer> Explanation: We can infer from the two sentences that the room type are both bedroom. Example 3: <Sentence1>: A small crowd gathered around the injured person. <Sentence2>: A crowd jumps up and down to the tunes played by an artist. <Condition>: The number of people <answer>yes(3)</answer> Explanation: While both sentences mention crowds, it is important and unclear how many people there are. Now, apply these steps to the following sentences: <Sentence1>: {sentence1} <Sentence2>: {sentence2} <Condition>: {condition}
A.3 Hyperparameter Configurations and Other Ablation Study
This section details the hyperparameter settings and provides empirical justification for selections not covered in the main paper. The default configurations used for our main results are summarized in Table 6.
Table 6: Default hyperparameter configurations for the main results.
HyperParameter Default Value
$\lambda_{1}:\lambda_{2}:\lambda_{3}$ 1:0.25:1
$\mu_{1}:\mu_{2}:\mu_{3}$ 1:1.5:1
$R_{\text{base}}$ 0.5
$G$ 8
NN 24
In addition to the ablation studies presented in the main paper, we conducted sensitivity analyses on three key hyperparameters: the binary reward weights in Stage I, generation multiplicity ($G$) and the pairwise base reward ($R_{\text{base}}$).
Sensitivity to Reward Weights in PoLi-RL Stage I.
In Stage I, the Binary Judgment Reward serves as an auxiliary signal to prevent the model from converging to “safe” median scores. We test the sensitivity to the weight of this component ($\lambda_{2}$). Results in Table 7 show that PoLi-RL is highly robust to variations in $\lambda_{2}$, consistently establishing a strong foundation for Stage II training.
Table 7: Sensitivity to Binary Reward Weight ($\lambda_{2}$) in PoLi-RL’s Stage I.
Method $\lambda_{1}$ (Pointwise) $\lambda_{2}$ (Binary) $\lambda_{3}$ (Format) Spearman $\uparrow$ Pearson $\uparrow$
PoLi-RL (Stage I) 1.0 0.1 0.1 44.94 44.93
1.0 0.25 0.1 44.77 44.45
1.0 0.5 0.1 44.67 44.42
1.0 1.0 0.1 44.76 44.92
Impact of Generation Multiplicity.
We investigate the trade-off between exploration diversity and computational overhead by varying $G\in\{4,8,12\}$. As shown in Table 8, while increasing $G$ theoretically aids RL exploration, empirical results indicate that $G=4$ offers insufficient exploration coverage: while the restricted search space may suffice for basic scoring, it fails to generate the diverse candidates necessary for learning the fine-grained semantic distinctions in Stage II. Conversely, scaling to $G=12$ leads to performance saturation and incurs higher computational overhead. These results validate our selection of $G=8$ as the optimal configuration for balancing performance and efficiency.
Table 8: Ablation on generation multiplicity ($G$) in PoLi-RL both stages.
$G$ Method Spearman $\uparrow$ Pearson $\uparrow$
4 PoLi-RL(Stage I) 44.55 44.49
4 PoLi-RL(Stage II) 47.47 47.45
8 PoLi-RL(Stage I) 44.77 44.45
8 PoLi-RL(Stage II) 48.18 48.27
12 PoLi-RL(Stage I) 45.51 45.61
12 PoLi-RL(Stage II) 47.98 48.06
Rationale for Pairwise Base Reward.
The pairwise reward (Eq. 7) incorporates a constant base reward $R_{base}$ to ensure a guaranteed positive signal when the model correctly predicts the ranking direction (ordinality), even if the exact score gap (cardinality) is imprecise. We compare the default $R_{base}=0.5$ against removing it ($0.0$) or overweighting it ($0.75$). Table 9 demonstrates that $0.5$ yields the best performance, justifying the need for a balanced reward structure.
Table 9: Sensitivity analysis of the pairwise base reward ($R_{\text{base}}$).
Method $R_{\text{base}}$ Spearman $\uparrow$ Pearson $\uparrow$
PoLi-RL (Stage II) 0 47.90 47.59
0.25 47.78 47.81
0.5 48.18 48.27
0.75 47.54 47.33
A.4 Comparison with Differentiable Ranking Objectives
In this section, we explicitly compare PoLi-RL against a strong cross-encoder regression baseline trained with differentiable ranking objectives. The goal is to determine whether optimizing a surrogate loss is sufficient to capture the rank-based nuances of C-STS. Specifically, we compare PoLi-RL against a baseline trained with the Pearson Correlation Coefficient (Pcc) Loss (Zhang & Li, 2024a), a state-of-the-art differentiable proxy for Spearman metric. Pcc loss is defined as:
$\mathcal{L}=1-\frac{cov(X,Y)}{\sigma_{X}\sigma_{Y}}$ (10)
where $X$ represents the predicted scores and $Y$ is the ground-truth labels. Table 10 shows that while the regression baseline outperforms standard SFT, PoLi-RL still maintains a clear advantage (1.74). This superiority stems from the paradigm shift: Regression treats the model as a ”black box” that maps embeddings to a scalar. In contrast, PoLi-RL optimizes the reasoning process itself. By aligning the token-level generation probability with the non-differentiable ranking metric, the model learns how to reason towards the correct score, enabling better generalization in complex conditional scenarios.
Table 10: Comparison between PoLi-RL and SFT with differentiable Pearson Correlation Coefficient (Pcc) Loss regression.
Methods Training Paradigm Spearman $\uparrow$ Pearson $\uparrow$
Qwen3-8B Few-shot 37.90 38.54
Qwen3-8B SFT (Auto-regressive) 40.42 40.83
Qwen3-8B SFT (Regression) 46.44 46.59
PoLi-RL (Qwen3-8B) RL 48.18 48.27
A.5 Performance on Re-annotated C-STS Dataset
Recent work by Tu et al. (2024) identified potential label noise in the original C-STS dataset and released a re-annotated validation set. Subsequently, Zhang et al. (2025c) further refined the dataset by correcting the condition descriptions and utilizing LLMs to clean the training labels. To ensure the robustness of our method against data quality issues, we re-evaluate PoLi-RL on the refined dataset provided by Zhang et al. (2025c).
As detailed in Table 11, PoLi-RL maintains its significant performance advantage on the re-annotated data, achieving a state-of-the-art Spearman correlation of 76.08 and outperforming the SFT baseline by nearly 4 points. This confirms that our reported improvements are robust and valid, rather than an artifact of overfitting to label noise.
Table 11: Performance on the Re-annotated C-STS Dataset. Evaluation performed on the 30% hold-out split of the validation set following the protocol of Tu et al. (2024) and Zhang et al. (2025c). Results marked with $\dagger$ are obtained from (Tu et al., 2024).
[…]
GPT-3.5 Few-shot - 66.1 64.8
GPT-4 Few-shot - 64.4 60.2
Qwen3-8B Few-shot 8B 64.42 64.5
Qwen3-8B SFT 8B 72.09 70.41
PoLi-RL (Qwen3-8B) RL 8B 76.08 74.16
A.6 LLM USAGE STATEMENT
The large language model (LLM) was utilized during the preparation of this manuscript. The use of this technology was strictly confined to the role of a writing assistant for the sole purpose of improving the linguistic quality of the text. Specifically, the LLM was employed for tasks related to grammar, syntax, phrasing, and overall readability. Its function was exclusively to perform surface-level linguistic refinements on text already written by the human authors. Crucially, the LLM did not contribute to any substantive or intellectual aspects of the research. The conceptualization of the study, the design of the methodology, the execution of experiments, the interpretation of results, and the formulation of conclusions were all executed by the human authors.